### v1

# Training Task Reasoning LLM Agents for Multi-turn Task Planning via Single-turn Reinforcement Learning

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in knowledge acquisition, reasoning, and tool use, making them promising candidates for autonomous agent applications. However, training LLM agents for complex multi-turn task planning faces significant challenges, including sparse episode-wise rewards, credit assignment across long horizons, and the computational overhead of reinforcement learning in multi-turn interaction settings. To this end, this paper introduces a novel approach that transforms multi-turn task planning into single-turn task reasoning problems, enabling efficient policy optimization through Group Relative Policy Optimization (GRPO) with dense and verifiable reward from expert trajectories. Our theoretical analysis shows that GRPO improvement on single-turn task reasoning results in higher multi-turn success probability under the minimal turns, as well as the generalization to subtasks with shorter horizons. Experimental evaluation on the complex task planning benchmark demonstrates that our 1.5B parameter model trained with single-turn GRPO achieves superior performance compared to larger baseline models up to 14B parameters, with success rates of 70% for long-horizon planning tasks with over 30 steps. We also theoretically and empirically validate the strong cross-task generalizability that the models trained on complex tasks can lead to the successful completion of all simpler subtasks.
1 Introduction
Large Language Models (LLMs) as autonomous agents are important in modern AI-based systems, which can perceive environments, reason about plans, and execute actions to interact with the environments [1]. Modern LLM agents demonstrate strong capabilities in knowledge integration, multi-step reasoning, and adaptive planning, as evidenced by their success in applications ranging from web search to robotic control [2, 3]. On top of these capabilities, prompt-based agentic frameworks [4, 5, 6] are proposed by integrating observation for environment state, reasoning based on augmented LLM with tools and memory, and action execution that interacts with the environment through structured interfaces as a series of single-turn interactions with the environments. However, effort-costly prompt engineering is inevitable to build the LLM-based agent, and it is also computationally expensive for test-time scaling in the multi-turn interaction with the environment [7, 8].
Therefore, training LLM agents through reinforcement learning (RL) for complex multi-turn task planning becomes a promising way to build effective agentic systems with low test-time cost [9, 10, 11]. However, current RL approaches face critical challenges when applied to multi-turn interactions with the environment for LLMs [12, 13, 14, 15]. The reward of task completion is too sparse for effective exploitation of LLM policy in the multi-turn settings. Besides, the credit assignment becomes particularly difficult in long-horizon planning problems, where determining which specific actions contribute to the eventual reward of success or failure is non-trivial [10]. The computational complexity of multi-turn RL scales poorly with sequence length, making it impractical for training on complex tasks that may require dozens of sequential decision turns [16].
In contrast, recent RL post-training successfully boosts LLM reasoning capability, particularly in math and coding domains based on on-policy methods, e.g., Group Relative Policy Optimization (GRPO) [17, 18]. However, these successes have primarily focused on single-turn reasoning tasks with verifiable reward, where the model generates a complete solution in one step and receives the reward directly based on the correctness of the final answer [19]. The fundamental challenge lies in bridging the gap between single-turn reasoning capabilities and multi-turn task planning requirements, where LLM agents must be coherent along the interaction sequences while adapting to changing environmental states.
To this end, we aim to bridge the gap based on the insight that complex multi-turn task planning can be decomposed into a sequence of single-turn task reasoning problems, each requiring the agent to select the optimal next action given the current state. This decomposition enables us to leverage the provably effective and efficient RL post-training with dense and verifiable reward based on expert trajectory, with theoretical guarantees of the multi-turn decision-making capabilities for long-horizon planning. We compare and validate the effectiveness and generalizability of the proposed method on the challenging long-horizon task planning benchmark. The contributions are summarized below.
  • We provide a theoretical analysis bridging single-turn task reasoning and multi-turn task completion, proving that single-turn GRPO improves the success probability in multi-turn planning scenarios.
  • We show empirical results on multiple tasks from the challenging Robotouille benchmark, where our 1.5B model trained with single-turn GRPO consistently outperforms 14B baselines with fewer completion steps.
  • We both theoretically and empirically show the strong cross-task generalization capabilities of the GRPO-optimized LLM agents, where models trained on complex tasks successfully transfer to simpler subtasks.
2 Problem Formulation
We formulate the multi-turn task planning problem as two interconnected Markov Decision Processes (MDPs): a multi-turn MDP representing the complete long-horizon task planning and a single-turn MDP based on expert trajectories for training purposes.
2.1 Multi-Turn Task Planning MDP
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},f,R,\mathcal{T},s_{0})$ be a finite-horizon MDP representing the multi-turn task planning problem, where $\mathcal{S}$ is the token state space representing the current task context and environment observation; $\mathcal{A}$ is the token action space of discrete possible single-step actions; $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the deterministic but unknown token state transition function; $R:\mathcal{S}\times\mathcal{A}\rightarrow\{0,1\}$ is the binary sparse reward function indicating whether the task is completed; $\mathcal{T}\subset\mathbb{N}$ is the finite time horizon; $s_{0}\in\mathcal{S}$ is the initial state. Let $\pi:\mathcal{S}\rightarrow\mathcal{A}$ be a stochastic policy mapping token states to token action distributions. $\pi(a|s)$ denotes the probability of $a=\pi(s)$. A trajectory $\tau=[(s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T},a_{T})]$ is generated by following policy $\pi$ for $T$ steps, where $s_{i+1}=f(s_{i},a_{i})$. Then we define the successful trajectory induced by a stochastic policy as follows.
Definition 2.1 (Successful trajectory by stochastic policy).
Given the MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},f,R,\mathcal{T},s_{0})$, a successful trajectory starting with $s_{0}^{s}=s_{0}$ is defined as:
$\tau^{s}(T)=[(s_{0}^{s},a_{0}^{s}),\ldots,(s_{T}^{s},a_{T}^{s})],\ R(s_{T}^{s},a_{T}^{s})=1,$ (1)
where $s_{i+1}^{s}=f(s_{i}^{s},a_{i}^{s})$ for $i=0,1,\dots,T-1$; $a_{i}^{s}\sim\pi(\cdot|s_{i}^{s})$ for $i=0,1,2,\dots,T$ given the stochastic policy $\pi$.
2.2 Expert Trajectory with Unique Optimality
On top of Definition˜2.1, we further define the expert trajectory generated by the expert policy with the unique optimality of minimal trajectory length.
Definition 2.2 (Expert trajectory by stochastic policy).
Given the MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},f,R,\mathcal{T},s_{0})$, a trajectory $\tau^{GT}(T^{GT})$ from the stochastic policy $\pi$ is defined as an expert trajectory if the following conditions hold:
  1. 1.
    $\tau^{GT}(T^{GT})$ is a successful trajectory generated from stochastic policy the $\pi$ in Definition˜2.1.
  2. 2.
    For any successful trajectory $\tau(T)$ generated by stochastic policy $\pi$ with $R(s_{T},a_{T})=1$, we have $T\geq T^{GT}$ and the equality holds if and only if $\tau=\tau^{GT}$.
Based on the expert trajectory $\tau^{GT}$ in Definition˜2.2, we can further define the expert policy $\pi^{GT}$ as follows.
Definition 2.3 (Expert policy from expert trajectory).
Given the MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},f,R,\mathcal{T},s_{0})$ and an expert trajectory $\tau^{GT}(T^{GT})$, a policy $\pi^{GT}:\mathcal{S}\rightarrow\mathcal{A}$ is defined as the expert policy if it holds that
$a_{i}^{GT}=\pi^{GT}(s_{i}^{GT}),\ \forall(s_{i}^{GT},a_{i}^{GT})\in\tau^{GT}(T^{GT}).$ (2)
Remark 2.1.
Even though expert trajectory $\tau^{GT}(T^{GT})$ is uniquely optimal with minimal length of turns $T^{GT}$, the induced expert policy $\pi^{GT}$ is not necessarily unique or deterministic, as there is no state-action constraint beyond expert trajectories $\tau^{GT}(T^{GT})$.
2.3 Single-Turn Task Reasoning MDP
Even though multi-turn task planning MDP $\mathcal{M}$ is the conventional one in the control and RL community, it is challenging and non-trivial to directly train an LLM policy for $\mathcal{M}$ due to intractable computational overhead with exponentially increasing token length of trajectory rollouts in the multi-turn setting [16, 10]. Therefore, to enable efficient LLM policy learning , we construct an additional single-turn MDP $\mathcal{M}_{S}=(\mathcal{S},\mathcal{A},\emptyset,r_{\pi^{GT}},\mathit{1},s_{0})$ based on the expert trajectory $\tau^{GT}(T^{GT})$ in Definition˜2.2 and expert policy $\pi^{GT}$ in Definition˜2.3. This single-turn MDP can be viewed as a bandit problem as follows. $\mathcal{S}$ is the same token state space as in $\mathcal{M}$, $\mathcal{A}$ is the same token action space as in $\mathcal{M}$, $r_{\pi^{GT}}:\mathcal{S}\times\mathcal{A}\rightarrow\{0,1\}$ is the binary reward function. There is no state transition function, which is denoted as $\emptyset$. The horizon $\mathit{1}$ is one-step in the single-turn setting. Initial state is the same as in $\mathcal{M}$.
Definition 2.4 (Reward function from expert trajectory).
The binary reward function $r_{\pi^{GT}}$ for single-turn MDP $\mathcal{M}_{S}=(\mathcal{S},\mathcal{A},\emptyset,r_{\pi^{GT}},\mathit{1},s^{GT})$ is defined as:
$r_{\pi^{GT}}(s,a)=\mathds{1}\{a=\pi^{GT}(s)\}\in\{0,1\},$ (3)
where $\pi^{GT}$ is the expert policy defined in Definition˜2.3.
The single-turn MDP $\mathcal{M}_{S}$ is based on expert trajectories $\tau^{GT}$, which enables efficient policy optimization with dense reward in Definition˜2.4. From the methodology section below, we will show that good performance on single-turn task reasoning (high reward $r$ in $\mathcal{M}_{S}$) will lead to successful long-horizon task planning (high reward $R$ in $\mathcal{M}$).
3 Methodology
We first give the theoretical analysis of policy learning for single-turn MDP $\mathcal{M}_{S}$ based on expert trajectories. Then we define the success probability with the minimal turns for task planning MDP $\mathcal{M}$ and show that policy improvement from $\mathcal{M}_{S}$ can lead to higher success probability in $\mathcal{M}$. Finally, we show that the learned policy can generalize well to unseen task planning with fewer completion turns.
3.1 GRPO for Single-Turn Task Reasoning
GRPO is an on-policy model-free RL method that optimizes policies using verifiable rewards and replaces critic model with a group of multiple responses for a prompt to calculate a group-based advantage [17, 18]. Based on the theoretical foundation of GRPO [19], we give the policy iteration and the success amplification in the single-turn MDP $\mathcal{M}_{S}$ defined with Definition˜2.4.
3.1.1 GRPO Policy Iteration Dynamics
Let $\rho_{Q}$ be a distribution over state query prompts $s\in\mathcal{S}\sim\rho_{Q}$. For the single-turn MDP $\mathcal{M}_{S}$, we consider the GRPO optimization problem which adopts the relative advantage within a group of policy outputs as the critic while maintaining proximity to a reference policy through KL regularization. The GRPO objective without clipping is formulated as:
$\max_{\pi}\mathbb{E}_{s\sim\rho_{Q}}\mathbb{E}_{a\sim\pi_{\text{old}}(\cdot|s)}\frac{\pi(a|s)}{\pi_{\text{old}}(a|s)}A(s,a)-\beta\text{KL}(\pi||\pi^{\text{ref}}),$ (4)
where the advantage function $A$ based on the binary reward in Definition˜2.4 is given by:
$A_{\pi^{GT}}(s,a)=\frac{r_{\pi^{GT}}(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\text{old}}(\cdot|s)}r_{\pi^{GT}}(s,a^{\prime})}{\sqrt{\text{Var}_{a^{\prime}\sim\pi_{\text{old}}(\cdot|s)}r_{\pi^{GT}}(s,a^{\prime})}}.$ (5)
At iteration $n\geq 1$, the policy iteration of GRPO for $\mathcal{M}_{S}$ is,
$\pi_{n}=\arg\max_{\pi}\mathbb{E}_{s\sim\rho_{Q}}\bigg\{\mathbb{E}_{a\sim\pi_{n-1}(\cdot|s)}\bigg[\frac{\pi_{n}(a|s)}{\pi_{n-1}(a|s)}\frac{r_{\pi^{GT}}(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{n-1}(\cdot|s)}r_{\pi^{GT}}(s,a^{\prime})}{\sqrt{\text{Var}_{a^{\prime}\sim\pi_{n-1}(\cdot|s)}r_{\pi^{GT}}(s,a^{\prime})}}\bigg]-\beta\text{KL}(\pi_{n}||\pi^{\text{ref}})\bigg\},$ (6)
where initial policy is the reference policy $\pi_{0}=\pi^{\text{ref}}$.
3.1.2 GRPO Single-Turn Optimality and Success Amplification
Following the Theorem˜3.1 from [19] below, we can have the optimality guarantee for GRPO iteration in Equation˜6 in our single-turn task reasoning setting.
Theorem 3.1 (Theorem 3 from [19]).
Let $\pi^{*}$ be the fixed point of GRPO-optimized policy obtained from the policy iteration dynamics of Equation˜6 and $\pi^{\text{ref}}$ be the initial reference policy. Denote the success probability of the reference policy as $p^{\text{ref}}=\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s^{GT})}[\mathds{1}\{a=\pi^{GT}(s^{GT})\}]$ and optimized policy as $p^{*}=\mathbb{E}_{a\sim\pi^{*}(\cdot|s^{GT})}[\mathds{1}\{a=\pi^{GT}(s^{GT})\}]$. It holds that $p^{*}>p^{\text{ref}}$.
Proof.
Proof can be found in Appendix D of [19]. ∎
Corollary 3.1 (GRPO single-Turn optimality).
For the fixed point of GRPO-optimized policy $\pi^{*}$ and the reference policy $\pi^{\text{ref}}$, the following holds for any state $s$ from the expert trajectory $\tau^{GT}$,
$\mathbb{E}_{a^{*}\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a^{*})]\geq\mathbb{E}_{a^{\text{ref}}\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a^{\text{ref}})].$ (7)
Proof.
It holds based on the single-turn reward formulation in Definition˜2.4 based on Theorem˜3.1. ∎
Corollary˜3.1 establishes that GRPO training on the single-turn MDP $\mathcal{M}_{S}$ with expert trajectories leads to policies that achieve higher success rates than the reference policy before GRPO training, providing the foundation for analyzing multi-turn performance improvements.
3.2 Success Probability for Multi-Turn Task Planning
To analyze how GRPO improvements on the single-turn MDP $\mathcal{M}_{S}$ can result in performance gains in the multi-turn MDP $\mathcal{M}$, we introduce the success probability with the minimal turns that captures the probability of achieving the optimal expert trajectory, by assuming the existence of minimal completion steps given any state below.
Definition 3.1 (Minimal turns for task completion).
For any state $s\in\mathcal{S}$ in the multi-turn MDP $\mathcal{M}$, denote the unique minimal length of turns as $T^{*}(s)\geq 0$ among all successful trajectories defined in Definition˜2.1 from $s$, i.e., $T^{*}(s):=\min\{T\geq 0:\exists\tau(T)\text{ s.t. }R(s_{T},a_{T})=1,s_{0}=s\}.$
$T^{*}(s)$ exists uniquely based on the existence and uniqueness of the expert trajectory in Definition˜2.2 by taking any states as initial states. We then define the probability of the successful trajectory with minimal steps.
Definition 3.2 (Success probability with minimal steps).
With $T^{*}(s)$ in Definition˜3.1, the probability of achieving the task planning goal in the minimal number of steps starting from state $s_{t}$ at time $t$ under policy $\pi$ in the multi-turn MDP $\mathcal{M}$ is defined as,
$P_{t}^{\pi}(s_{t})=\mathbb{P}_{\pi}\left(R(s_{t+T^{*}(s_{t})},a_{t+T^{*}(s_{t})})=1|s_{t}\right).$ (8)
We then present the recursion condition for the success probability in Definition˜3.2 based on the reward function from the expert trajectory in Definition˜2.4.
Theorem 3.2 (Recursion equation of success probability).
The success probability defined in Definition˜3.2 satisfies the following recursion condition,
$P_{t}^{\pi}(s_{t})=\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r_{\pi^{GT}}(s_{t},a)\cdot P_{t+1}^{\pi}(s_{t+1})],$ (9)
where $\pi^{GT}$ is the expert policy in Definition˜2.3 by expert trajectory with $P_{t+T^{*}(s_{t})}^{\pi}(s_{t+T^{*}(s_{t})})=1$ for any state $s_{t}$.
Proof.
For any state $s_{t}$ as the initial state, based on Definition˜2.3, $\pi^{GT}$ is induced by the expert trajectory $\tau^{GT}=[(s_{t},a_{t}),\dots,(s_{T^{GT}},a_{T^{GT}})],s_{k+1}=f(s_{k},\pi^{GT}(s_{k})),k=t,t+1,\dots,T^{GT}-1$ with minimal turns of task completion. By $P_{t+T^{*}(s_{t})}^{\pi}(s_{t+T^{*}(s_{t})})=1$, we have $R(s_{t+T^{*}(s_{t})},a_{t+T^{*}(s_{t})})=1$. Combining the uniqueness of expert trajectory in Definition˜2.2 and minimal completion steps $T^{*}(s)$ in Definition˜3.1, it holds that
$T^{*}(s_{T^{GT}})=0,T^{GT}=T^{*}(s_{k})+k,k=t,\dots,T^{GT}.$ (10)
For any state that deviates from the expert trajectory $s_{k+1}^{\prime}\neq f(s_{k},\pi^{GT}(s_{k}))$, by the minimal completion turns,
$T^{*}(s_{k+1}^{\prime})>T^{*}(s_{k})-1,\forall k=t,t+1,\dots,T^{GT}-1.$ (11)
By the law of total expectation over the next action $t+1$:
$P_{t}^{\pi}(s_{t})=\mathbb{P}_{\pi}(R(s_{t+T^{*}(s_{t})},a_{t+T^{*}(s_{t})})=1|s_{t})$ (12)
$=\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[\mathbb{P}_{\pi}(R(s_{t+T^{*}(s_{t})},a_{t+T^{*}(s_{t})})=1|s_{t},a)].$
For $a\sim\pi(\cdot|s_{t})$, if $a\neq\pi^{GT}(s_{t}),i.e.,r_{\pi^{GT}}(s_{t},a)=0$, we have $s_{t+1}^{\prime}\neq f(s_{t},\pi(s_{t}))$. Then based on Equation˜11, it takes strictly more than $T^{*}(s_{t})-1$ steps for completion, i.e.,
$\mathbb{P}_{\pi}(R(s_{t+1+T^{*}(s_{t})-1},a_{t+1+T^{*}(s_{t})-1})=1|s_{t+1}^{\prime})=0.$ (13)
If $a=\pi^{GT}(s_{t}),i.e.,r_{\pi^{GT}}(s_{t},a)=1,s_{t+1}=f(s_{t},a)$, then based on Equation˜10, we have
$\mathbb{P}_{\pi}(R(s_{t+T^{*}(s_{t})},a_{t+T^{*}(s_{t})})=1|s_{t},a)$ (14)
$=$ $\mathbb{P}_{\pi}(R(s_{t+1+T^{*}(s_{t+1})},a_{t+1+T^{*}(s_{t+1})})=1|s_{t+1})$
$=$ $P_{t+1}^{\pi}(s_{t+1}).$
With the values of $r_{\pi^{GT}}(s_{t},a)\in\{0,1\}$ and combining Equations˜12, 13 and 14, it holds that
$P_{t}^{\pi}(s_{t})=\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r_{\pi^{GT}}(s_{t},a)\cdot P_{t+1}^{\pi}(s_{t+1})],$ (15)
which concludes the proof. ∎
Please note that the recursion equation in Theorem˜3.2 is not the same as Bellman recursion in value function iteration [20], and the key difference lies in that the multiplication of the binary single-turn reward $r_{\pi^{GT}}(s_{t},a)$ is involved in the recursion of expectation, instead of addition. This recursion bridges the final multi-turn success in Equation˜8 with single-turn reward improvements at each step in Equation˜7, which is crucial to the following analysis on how single-turn GRPO improves the multi-turn task planning.
3.3 GRPO Improvement of Multi-Turn Success Probability
We now analyze how GRPO improvements on single-turn task reasoning in $\mathcal{M}_{S}$ will lead to performance gains in the multi-turn task planning MDP $\mathcal{M}$. First, we assume the GRPO-optimized policy for single-turn task reasoning based on expert trajectories can generalize to all states.
Assumption 3.1 (GRPO improvement on all states).
The GRPO improvement on expert trajectory states in Equation˜7 also holds for all states $s\in\mathcal{S}$, i.e., $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)]\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a)],$ where $\pi^{*}$ is the GRPO-optimized policy from Corollary˜3.1.
The assumption above is a slight relaxation of Corollary˜3.1, which empirically holds when the policy model is fine-tuned from large language models (LLMs) with strong generalizability [21, 22]. We then present how the GRPO will improve the multi-turn task planning with a higher success probability with the fewest turns.
Theorem 3.3 (GRPO improves success probability).
Let $\pi^{*}$ be the fixed point of the GRPO-optimized policy from the policy iteration dynamics of Equation˜6 and $\pi^{\text{ref}}$ be the initial reference policy. Under Assumption 3.1 and Equation˜7, it holds for any given state $s_{t}$ at step $t$ that,
$P_{t}^{\pi^{*}}(s_{t})\geq P_{t}^{\pi^{\text{ref}}}(s_{t})\quad\forall s_{t}\in\mathcal{S}.$ (16)
Proof.
By the existence and uniqueness of the expert trajectory in Definition˜2.2, denote the expert trajectory starting from $s_{t}$ as $\tau^{GT}=[(s_{t},a_{t}),\dots,(s_{T^{GT}},a_{T^{GT}})]$ with minimal turns of task completion $T^{*}(s_{t})=T^{GT}-t$ in Definition˜3.1. We prove by backward induction on the length of remaining turns to task completion. In the final case when $t=T^{GT}$, we have $P_{t}^{\pi^{*}}(s_{t})=P_{t}^{\pi^{\text{ref}}}(s_{t})=1$ since the task is already completed. Now for the inductive case at step $t+1,t+1\leq T^{GT}$, we assume $P_{t+1}^{\pi^{*}}(s^{\prime})\geq P_{t+1}^{\pi^{\text{ref}}}(s^{\prime})$ for all states $s^{\prime}$. Then for any state $s_{t}$ at step $t$, it holds that
$P_{t}^{\pi^{*}}(s)$ $=\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{*}}(f(s,a))]$ (17)
$\geq\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]$ (18)
$\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]$ (19)
$=P_{t}^{\pi^{\text{ref}}}(s),$ (20)
where Equations˜17 and 20 are from Theorem˜3.2, Equation˜18 is based on the inductive hypothesis at step $t+1$ and Equation˜19 is based on Assumption 3.1. ∎
Furthermore, we show that the policy trained with expert trajectories with single-turn GRPO can also generalize to subtask defined below.
Definition 3.3 (Subtask and its success probability).
Given the expert trajectory $\tau^{GT}(T^{GT})=[(s_{0},a_{0}),\dots,(s_{T^{GT}},a_{T^{GT}})]$ in Definition˜2.2 for the full task, a subtask is defined by a completion step $k^{*}<T^{GT}$ where the subtask is completed upon reaching $(s_{k^{*}},a_{k^{*}})$, i.e., $R_{\text{sub}}(s_{k^{*}},a_{k^{*}})=1$. The success probability with the fewest turns for subtask at $s_{t},t\leq k^{*}$ is defined as,
$P_{t}^{\text{sub},\pi}(s_{t})=\mathbb{P}_{\pi}\left(R_{\text{sub}}(s_{k^{*}},a_{k^{*}})=1|s_{t}\right),$ (21)
which is the probability of achieving subtask goal in the minimal number of steps starting from state $s_{t}$ at time $t$ under policy $\pi$.
Based on the success probability of subtask, we present the following corollary of Theorem˜3.3.
Corollary 3.2 (GRPO generalization to subtasks).
For any subtask defined by completion step $k^{*}<T^{GT}$ in Definition˜3.3, the GRPO-optimized policy $\pi^{*}$ achieves better performance on the subtask than the reference policy $\pi^{\text{ref}}$ for any given state $s_{t}$ at step $t\leq k^{*}$, i.e., $P_{t}^{\text{sub},\pi^{*}}(s_{t})\geq P_{t}^{\text{sub},\pi^{\text{ref}}}(s_{t})$.
Proof.
Since the subtask uses the same expert trajectory prefix as the full task, Assumption 3.1 also holds for subtask goal with the same single-turn reward function $r_{\pi^{GT}}(s,a)$. Besides, the recursion condition of $P_{t}^{\text{sub},\pi}(s_{t})$ also applies. Therefore, following the proof of Theorem˜3.3, by letting $T^{GT}=k^{*}$, $\pi^{*}$ achieves better success probability than $\pi^{\text{ref}}$ over $P_{t}^{\text{sub}}(s_{t})$ for any given state $s_{t}$ at step $t\leq k^{*}$. ∎
4 Experiments
In the experiments, we aim to answer the following two questions empirically: 1) How is the performance of multi-turn task planning after GRPO over single-turn task reasoning? 2) Can the LLM agents after GRPO generalize to unseen planning tasks? We answer the first question in Section˜4.2 and the second one in Section˜4.3, which correspond to Theorem˜3.3 and Corollary˜3.2, respectively. Prior to the answers, we first introduce the experimental setup of environment, model training and evaluation metrics.
4.1 Experimental Setup
Environment and Tasks
We evaluate our approach on the Robotouille benchmark [23], a challenging multi-turn task planning environment designed for testing LLM agents’ reasoning capabilities in cooking scenarios. Robotouille provides a deterministic long-horizon complex environment where agents must coordinate multiple actions across different kitchen stations to complete recipes. We focus on four progressively challenging cooking tasks with different maximum steps: (1) Cheese Sandwich - at most 10 steps; (2) Burger - at most 10 steps; (3) Cheese Burger - at most 15 steps; and (4) Double Cheese Burger - at most 23 steps. These tasks were selected to provide increasing complexity in both planning horizon and reasoning requirements, allowing us to evaluate our method’s scalability and generalization across different task difficulties. While Robotouille represents a specific domain, its structured nature with consistent state-action distribution under varying complexity levels makes it particularly suitable as a controlled yet comprehensive testbed to empirically validate our theoretical results. Future work will extend it to more diverse agent environments [24, 25, 15] to demonstrate the broader applicability of single-turn decomposition for multi-turn agent training.
Table 1: Performance Comparison on Different Tasks. Best results are in bold. “—” indicates no successful trajectories.
Results of Cheese Sandwich SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
Llama3.3-70b (expert) 0.5 11.9 8.8
Qwen2.5-1.5b 0.0 15.0
Qwen2.5-3b 0.0 15.0
Qwen2.5-7b 0.0 15.0
Qwen2.5-14b 0.2 13.6 8.0
Qwen2.5-1.5b (w/ SFT) 0.0 15.0
Qwen2.5-1.5b (w/ SFT & GRPO) 0.3 12.5 6.7
Results of Burger SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
Llama3.3-70b (expert) 0.4 12.6 9.0
Qwen2.5-1.5b 0.0 15.0
Qwen2.5-3b 0.0 15.0
Qwen2.5-7b 0.1 14.9 14.0
Qwen2.5-14b 0.0 15.0 15.0
Qwen2.5-1.5b (w/ SFT) 0.2 13.7 8.5
Qwen2.5-1.5b (w/ SFT & GRPO) 0.7 10.1 8.0
Results of Cheese Burger SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
Llama3.3-70b (expert) 0.7 17.4 15.0
Qwen2.5-1.5b 0.0 23.0
Qwen2.5-3b 0.0 23.0
Qwen2.5-7b 0.0 23.0
Qwen2.5-14b 0.2 22.4 20.0
Qwen2.5-1.5b (w/ SFT) 0.6 18.4 15.3
Qwen2.5-1.5b (w/ SFT & GRPO) 0.7 15.8 12.7
Results of Double Cheese Burger SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
Llama3.3-70b (expert) 0.1 33.5 20.0
Qwen2.5-1.5b 0.0 35.0
Qwen2.5-3b 0.0 35.0
Qwen2.5-7b 0.0 35.0
Qwen2.5-14b 0.0 35.0
Qwen2.5-1.5b (w/ SFT) 0.1 34.2 27.0
Qwen2.5-1.5b (w/ SFT & GRPO) 0.3 30.5 20.0
Table 2: Cross-task Generalization of Success Rate (SR)
Evaluated ($\rightarrow$)
Trained ($\downarrow$)
Cheese
Sandwich
Burger
Cheese
Burger
Double Cheese
Burger
Cheese Sandwich 0.3 0.3 0.0 0.0
Burger 0.4 0.7 0.0 0.0
Cheese Burger 0.4 0.7 0.7 0.0
Double Cheese Burger 0.3 0.5 0.4 0.3
Expert Trajectory Collection
Expert trajectories are collected using rejection sampling from Llama3.3-70B-Instruct [21], which serves as our expert policy $\pi^{GT}$. To ensure diversity and robustness, we generate expert trajectories across varied kitchen layouts and ingredient placements, creating approximately 100 successful trajectories per task type. Each trajectory $\tau^{GT}=[(s_{0}^{GT},a_{0}^{GT}),\ldots,(s_{n}^{GT},a_{n}^{GT})]$ satisfies the task completion criteria with $R(s_{n}^{GT},a_{n}^{GT})=1$. The state-action tokens follow the ReAct agentic framework [4], where each action $a_{i}^{GT}$ consists of structured reasoning followed by a specific action selection from the valid action list indicated from state tokens. This framework enables us to construct single-turn verifiable rewards $r(s_{i}^{GT},\pi_{\theta}(s_{i}^{GT}))=\mathds{1}\{\pi_{\theta}(s_{i}^{GT})=a_{i}^{GT}\}$ for GRPO training, where both the ReAct format and action correctness contribute to the reward by 0.1 and 0.9, respectively.
Model Training Pipeline
Our training pipeline consists of two sequential post-training stages using the VeRL implementation [26] of GRPO. First, we perform supervised fine-tuning (SFT) on Qwen2.5-1.5B-Instruct [22] using the collected expert trajectories to establish a strong initialization. The SFT stage trains the model to imitate the ReAct pattern for the action tokens. Subsequently, we apply GRPO with the single-turn verifiable reward function based on all state-action pairs $(s_{i}^{GT},a_{i}^{GT})$ from expert trajectories. The SFT takes 8 epochs with batch size of 8 and the GRPO training uses a batch size of 256, learning rate of $1\times 10^{-6}$, and KL regularization coefficient $\beta=0.001$ for 50 epochs.
Table 3: Cross-task Generalization of Average Steps of All Trajectories (ASAT)
Evaluated ($\rightarrow$)
Trained ($\downarrow$)
Cheese
Sandwich
Burger
Cheese
Burger
Double Cheese
Burger
Cheese Sandwich 12.5 13.1 23 35
Burger 13.5 10.1 23 35
Cheese Burger 12.2 10.8 15.8 35
Double Cheese Burger 13.3 11.6 19.4 30.5
Table 4: Cross-task Generalization of Average Steps of Successful Trajectories (ASST). “—” indicates no successful trajectories.
Evaluated ($\rightarrow$)
Trained ($\downarrow$)
Cheese
Sandwich
Burger
Cheese
Burger
Double Cheese
Burger
Cheese Sandwich 6.7 8.67
Burger 11.25 8.0
Cheese Burger 8.0 9.0 12.7
Double Cheese Burger 9.3 8.2 14 20
Evaluation Metrics and Baselines
We evaluate all models on 10 unseen kitchen layouts per task type, ensuring that the test environments differ from training data in terms of object positions, kitchen topology, and ingredient availability. The timeout lengths of attempt turns are 15, 15, 23, 35 for tasks of Cheese Sandwich, Burger, Cheese Burger, Double Cheese Burger, respectively. We assess the model performance of effectiveness and efficiency using three key metrics: (1) Success Rate (SR) - the proportion of episodes where the agent successfully completes the task within the step timeout limit; (2) Average Steps of All Trajectories (ASAT) - the mean number of steps taken across all episodes, including timeout failures; and (3) Average Steps of Successful Trajectories (ASST) - the mean number of steps for successfully completed episodes only. We compare our approach against different-size (1.5B, 3B, 7B, 14B) Qwen2.5 instruct models [22] using ReAct prompting [4] and the one right after SFT on expert trajectories but without GRPO training, ablating the contribution of RL for task reasoning.
4.2 Results of Task Planning with Single-turn GRPO
Table˜1 presents the results for the four progressively challenging cooking tasks. Each table compares the proposed method against multiple baselines including different-sized Qwen models and the model without single-turn task reasoning of GRPO. These results provide strong empirical support for Theorem˜3.3, showing that single-turn GRPO training indeed leads to improved multi-turn task planning performance. Specifically, the results demonstrate the following key findings.
Significant Performance Gains by GRPO of Task Reasoning
Our model with SFT and GRPO consistently outperforms much larger-size baseline models. Most notably, it achieves 0.70 success rate on the Burger task compared to 0.0 for the 14B baseline, and matches the expert’s 0.70 performance on the Cheese Burger task while being significantly more efficient (15.8 vs 17.4 average steps). The SFT-alone model provides limited success compared to the one followed by GRPO, validating the effectiveness of single-turn RL for task reasoning in Theorem˜3.3.
Efficiency and Scalability to Complex Tasks
Beyond success rates, our method consistently achieves lower average step counts for successful episodes, indicating more efficient task planning. For instance, on the Cheese Burger task, our method requires only 12.7 steps compared to 15.3 for SFT-only, demonstrating that GRPO learns more optimal policies with minimal steps of task completion in Definition˜3.2. The performance advantage is maintained even on the most challenging Double Cheese Burger task (35-step horizon), where our 1.5B model achieves 0.3 success rate while all larger LLMs fail completely.
4.3 Generalization to Unseen Task Planning
To evaluate the generalizability of LLM agent trained from single-turn task reasoning, we train LLM agents for individual tasks and evaluate their zero-shot performance on all four tasks. Tables 2, 3, and 4 present the results under each metric. These findings provide empirical validation of Corollary˜3.2 and demonstrate that our single-turn GRPO approach enables strong cross-task generalization, particularly in the direction from complex to simpler tasks, which is valuable for practical deployment scenarios where a single model needs to handle unseen varying task complexities. Some specific findings are shown below.
Consistent Generalization of Success Rate over Subtasks
Models trained on more complex tasks demonstrate better generalization to simpler tasks. The agent trained on Double Cheese Burger can solve all tasks with non-zero success rates (0.3 - 0.5), while agents trained on simpler tasks fail on more complex ones. This supports our theoretical result that complex task training captures the necessary subtask planning capabilities in Corollary˜3.2, showing that GRPO training enables generalization to subtasks with shorter horizons. We also observe that success rates generally decrease as the evaluation task becomes more complex relative to the training task, but remain positive when the evaluation task is simpler or of similar complexity.
Loss of Optimality and Failure on Complex Tasks in Generalization
Even when agents trained on complex tasks successfully generalize to simple tasks, they may require more steps than the agents trained specifically on the target simple tasks. For example, the Double Cheese Burger-trained agent requires 8.2 steps for Burger task completion compared to 8.0 steps for the Burger-trained agent, indicating some loss of efficiency and optimality during generalization. In addition, agents trained on all simpler tasks completely fail on the most complex Double Cheese Burger task, revealing the future directions of the LLM RL post-training community where complex reasoning capabilities are hard to obtain from simpler task reasoning problems.
5 Conclusion
In this paper, we present a theoretical framework demonstrating that GRPO improvements on single-turn task reasoning result in enhanced multi-turn success probability under minimal steps for challenging multi-turn interaction scenarios. Empirical results validate our approach, with our small-parameter model consistently outperforming larger baselines while requiring fewer steps for task completion. We also show cross-task generalization capabilities both theoretically and empirically. While limitations remain regarding the reliance on expert trajectories, our work shows how to leverage single-turn RL successes for complex multi-turn task planning, opening promising directions for more capable and efficient LLM-based agents.